# Copula Density Neural Estimation

### v2

@@ III-A Variational Formulation
Notice that the density of the copula can be derived with the same approach also by working in the sample domain. Indeed, when $p$ and $q$ correspond to the joint and the product of the marginals, respectively, the following corollary holds.
Corollary 1.1.
then
is the copula density, where
and
$\hat{T}(\mathbf{x})=\arg\max_{T}\mathcal{J}_{f,x}(T),$ (14)
@@
where $c$ is the density of the copula obtained as in (1). The thesis then follows immediately from Theorem 1. ∎
A great advantage of the formulation in (5) comes from the second expectation term. Conversely to the variational discriminative formulation in (11) that tests jointly with marginals samples, the comparison in (5) is made between samples from the joint copula structure and independent uniforms. The latter can be easily generated without the need of any scrambler that factorizes $p_{X}$ into the product of the marginal pdfs. More precisely, a derangement type of shuffling mechanism would be required as proved in [15]. On the other hand, (5) needs samples from the copula, thus, needs an estimate of the marginals of $X$ to apply transform sampling. In the following, we use the formulation in (5) which inherently possesses another desired property when using neural networks.
III-B Parametric Implementation
$\hat{\theta}=\arg\max_{\theta}\mathcal{J}_{f}(T_{\theta}).$ (16)
Since at convergence the network outputs a transformation of the copula density evaluated at the input $\mathbf{u}$, the final layer possesses a unique neuron with activation function that depends on the generator $f$ (see the code for more details).
Name Generator $f(u)$ Conjugate $f^{*}(t)$
GAN $u\log u-(u+1)\log(u+1)+\log(4)$ $-\log(1-\exp(t))$
KL $u\log u$ $\exp(t-1)$
HD $(\sqrt{u}-1)^{2}$ $t/(1-t)$
The resulting estimator of the copula density reads as follows
and its training procedure enjoys two normalization properties. The former consists in a natural normalization of the input data in the interval $[0,1)$ via transform sampling that facilitates the training convergence and helps producing improved dependence measures [16]. The latter normalization property is perhaps at the core of the proposed methodology. The typical problem in creating neural density estimators is to enforce the network to return densities that integrate to one
IV Evaluation Measures
Figure 1: Approximation quality of $c_{V}$. Comparison between CODINE method $\hat{c}_{V}$ and the flat copula density $\pi_{V}$ for different values of the signal-to-noise ratio (SNR) and for different dimensionality $d$ of the input.
To assess the quality of the copula density estimator $\hat{c}_{U}(\mathbf{u})$, we propose the following set of self-consistency tests over the basic property illustrated in (18). In particular,
  1. 1.
  2. 2.
The first test verifies that the copula density integrates to one while the second set of tests extends the first test to the moments of any order. Similarly, joint consistency tests can be defined, e.g., the Spearman rank correlation $\rho_{X,Y}$ between pairs of variables can be rewritten in terms of their joint copula density $\hat{c}_{UV}$ and it reads as follows
When the copula density is known, it is possible to assess the quality of the copula density neural estimator $\hat{c}_{U}(\mathbf{u})$ by computing the Kullback-Leibler (KL) divergence between the true and the estimated copulas
Now, as a first example to validate the density estimator, we consider the transmission of $d$-dimensional Gaussian samples over an additive white Gaussian channel (AWGN). Given the AWGN model $Y=X+N$, where $X\sim\mathcal{N}(0,\mathbb{I})$ and $N\sim\mathcal{N}(0,\Sigma_{N})$, it is simple to obtain closed-form expressions for the probability densities involved. In particular, the copula density of the output $Y$ reads as in (IV), where $\mathbf{F}_{X}(\mathbf{x})$ is an operator that element-wise applies transform sampling (via Gaussian cumulative distributions) to the components of $\mathbf{x}$ such that $(\mathbf{u},\mathbf{v})=(\mathbf{F}_{X}(\mathbf{x}),\mathbf{F}_{Y}(\mathbf{y}))$ and $\tilde{\Sigma}_{N}=\Sigma_{N}\odot\mathbb{I}$, where $A\odot B$ denotes the Hadamard product between $A$ and $B$.
In Fig. 1 we illustrate the KL divergence (in bits) between the ground-truth and the neural estimator obtained using the $f$-GAN generator reported in Tab. I. To work with non-uniform copula structures, we study the case of non-diagonal noise covariance matrix $\Sigma_{N}$. In particular, we impose a tridiagonal covariance matrix such that $\Sigma_{N}=\sigma_{N}^{2}R$ where $R_{i,i}=1$ with $i=1,\dots,d$, and $R_{i,i+1}=\rho$, with $i=1,\dots,d-1$ and $\rho=0.5$. Moreover, Fig. 1 also depicts the quality of the approximation for different values of the signal-to-noise ratio (SNR), defined as the reciprocal of the noise power $\sigma_{N}^{2}$, and for different dimensions $d$. To provide a numerical comparison, we also report the KL divergence $D_{\text{KL}}(c_{V}||\pi_{V})$ between the ground-truth and the flat copula density $\pi_{V}=1$. It can be shown that when $c_{V}$ is Gaussian, we obtain
V Results
V-A Mutual Information Estimation
${I}(X;Y)=\mathbb{E}_{(\mathbf{u},\mathbf{v})\sim c_{UV}(\mathbf{u},\mathbf{v})}\biggl{[}\log\frac{c_{UV}(\mathbf{u},\mathbf{v})}{c_{U}(\mathbf{u})c_{V}(\mathbf{v})}\biggr{]}.$ (25)
Therefore, (25) requires three separate copula densities estimators, each of which obtained as explained in Section III. Alternatively, one could learn the copulas density ratio via maximization of the variational lower bound on the mutual information. Using again Fenchel duality, the KL divergence
corresponds to the supremum over $T$ of
$\small\mathcal{J}_{\text{KL}}(T)=\mathbb{E}_{c_{UV}}\biggl{[}T\bigl{(}\mathbf{u},\mathbf{v}\bigr{)}\biggr{]}-\mathbb{E}_{c_{U}c_{V}}\biggl{[}\exp{\bigl{(}T\bigl{(}\mathbf{u},\mathbf{v}\bigr{)}-1\bigr{)}}\biggr{]}.$ (27)
When $X$ is a univariate random variable, its copula density $c_{U}$ is unitary. Notice that (27) can be seen as a special case of the more general (5) when $f$ is the generator of the KL divergence and the second expectation is not done over independent uniforms with distribution $\pi_{U}$ but over samples from the product of copula densities $c_{U}\cdot c_{V}$.
We estimate the mutual information between $X$ and $Y$ in the AWGN model using (27) and the generators described in Tab. I. Fig. 3a and Fig. 3b show the estimated mutual information for $d=1$ and $d=5$, respectively, and compare it with the closed-form capacity formula $I(X;Y)=d/2\log_{2}(1+\text{SNR})$.
V-B Data Generation
As a second application example, we generate $n$ new pseudo-observations $\{\hat{\mathbf{u}}^{(1)},\dots,\hat{\mathbf{u}}^{(n)}\}$ from $\hat{c}_{U}$ by deploying a Markov chain Monte Carlo (MCMC) algorithm. Validating the quality of the generated data provides an alternative path for assessing the copula estimate itself.
V-B1 2D toy dataset
V-B2 MNIST dataset
As a more complex example, we report the generation of the MNIST handwritten digits of size $28\times 28$. In particular, we study the copula density of the latent space obtained from an autoencoder. With such experiment, similarly to the experiment in [13], we firstly prove that the approach works even for data of dimension $d=25$ and we secondly want to emphasize the fact that the approach is potentially scalable to higher dimensions. The idea is to train a convolutional neural network autoencoder to reconstruct digits (see Fig. 6). During training, the autoencoder learns latent representations (via an encoder mapping) that can be analyzed and synthesized with CODINE.
The analysis and synthesis strategies in Fig. 4 comprise a first marginal block to project the latent representations into the uniform probability space. Then, a CODINE block learns the copula density which is used by Gibbs sampling to generate new uniform latent representations. An inverse transform sampling block maps data from the uniform to the sample space. Once new latent samples are generated, it is possible to feed them into the pre-trained decoder and obtain new digits, as illustrated in Fig. 7. Although the CODINE block $T(\mathbf{u})$ is a simple shallow neural network, the generated digits visually resemble the original ones, meaning that our approach managed to, at least partially, estimate the density of the uniform latent representations.
Figure 6: Autoencoder architecture used to learn latent vectors ($d=25$). The autoencoder is trained with binary cross-entropy
V-C Generative model based on the copula estimate
@@
where $\mathcal{G}$ is a class of functions $g:\chi\rightarrow\mathbb{R}$. Since $c_{U}=c_{V}$ if and only if $\mathbb{E}_{\mathbf{u}\sim c_{U}(\mathbf{u})}[g(\mathbf{u})]=\mathbb{E}_{\mathbf{v}\sim c_{V}(\mathbf{v})}[g(\mathbf{v})]$ $\forall g\in\mathcal{G}$, MMD is a metric that measures the disparity between $c_{U}$ and $c_{V}$ (see [19]). If $g(\mathbf{u})=c_{U}(\mathbf{u})$ is a valid function, we can define a plausible loss function based on the MMD metric as follows
$\min_{\theta_{G}}\mathbb{E}_{\mathbf{u}\sim c_{U}(\mathbf{u})}[c_{U}(\mathbf{u})]-\mathbb{E}_{\mathbf{v}\sim\pi_{V}(\mathbf{v})}[c_{U}(G(\mathbf{v};\theta_{G}))].$ (29)
@@
$W(c_{U},c_{V})=\frac{1}{K}\sup_{||h||_{L}\leq K}\biggl{[}\mathbb{E}_{\mathbf{u}\sim c_{U}(\mathbf{u})}[h(\mathbf{u})]-\mathbb{E}_{\mathbf{v}\sim c_{V}(\mathbf{v})}[h(\mathbf{v}))\biggr{]}.$ (30)
Under such conditions, (29) can be interpreted as the generator loss function of a Wasserstein-GAN [20] where the optimum discriminator $h$ is supposed to be known and corresponds to the learnt copula density $c_{U}$. The proposed idea lies in between two established approaches. The first one, generative moment matching networks GMMNs [21], assume $\mathcal{G}$ as the reproducing kernel Hilbert space where $g$ is a kernel $k\in\mathcal{H}$ and the supremum in (28) is thus attained. Such MMD-based optimizes only over the generator’s parameters but does not produce expressive generators, mainly because of the restriction imposed by the kernel structure. The second, instead, requires to learn both the generator and the discriminator, the latter in order to reach the supremum in (30). However, enforcing the Lipschitz constraint is not trivial and the alternation between generator and discriminator training suffers from the usual instability and slow convergence problems [5]. Even if the copula-based approach does not claim optimality, it possesses two desirable properties: compared to kernel-based methods, it uses a more powerful and appropriate discriminator, the copula density itself. Moreover, the fact that $c_{U}$ is obtained from a prior analysis renders the generator learning process uncoupled from the discriminator’s one.
Figure 7: 100 randomly selected digits obtained using CODINE to generate the latent vector fed in the decoder. Comparison with test data.
VI Conclusions
This brief presented CODINE, a copula density neural estimator. It works by maximizing a variational lower bound on the $f$-divergence between two distributions defined on the uniform probability space, namely, the distribution of the pseudo-observations and the distribution of independent uniforms. CODINE also provides alternative approaches for measuring statistical dependence, mutual information, and for data generation.

### v3

@@ A-A Proof of Theorem 1
A-B Proof of Corollary 1.1
@@
where $c$ is the density of the copula obtained as in (1). The result then follows immediately from Theorem 1. ∎
A-C Generative Model Based on the Copula Estimate
@@
where $\mathcal{G}$ is a class of functions $g:\chi\rightarrow\mathbb{R}$. Since $c_{U}=c_{V}$ if and only if $\mathbb{E}_{\mathbf{u}\sim c_{U}(\mathbf{u})}[g(\mathbf{u})]=\mathbb{E}_{\mathbf{v}\sim c_{V}(\mathbf{v})}[g(\mathbf{v})]$ $\forall g\in\mathcal{G}$, MMD is a metric that measures the disparity between $c_{U}$ and $c_{V}$ (see [42]). If $g(\mathbf{u})=c_{U}(\mathbf{u})$ is a valid function, we can define a plausible loss function based on the MMD metric (and referred to as $\mathcal{J}_{\text{MMD}}(G_{\theta_{G}})$) as follows
$\min_{\theta_{G}}\mathbb{E}_{\mathbf{u}\sim c_{U}(\mathbf{u})}[c_{U}(\mathbf{u})]-\mathbb{E}_{\mathbf{v}\sim\pi_{V}(\mathbf{v})}[c_{U}(G_{\theta_{G}}(\mathbf{v}))].$ (31)
@@
$W(c_{U},c_{V})=\frac{1}{K}\sup_{||h||_{L}\leq K}\biggl{[}\mathbb{E}_{\mathbf{u}\sim c_{U}(\mathbf{u})}[h(\mathbf{u})]-\mathbb{E}_{\mathbf{v}\sim c_{V}(\mathbf{v})}[h(\mathbf{v}))\biggr{]},$ (32)
where $||\cdot||_{L}$ is the $L$-th norm. Under such conditions, (31) can be interpreted as the generator loss function of a Wasserstein-GAN [43] where the optimum discriminator $h$ is supposed to be known and corresponds to the learnt copula density $c_{U}$. The proposed idea lies in between two established approaches. The first one, generative moment matching networks GMMNs [44], assume $\mathcal{G}$ as the reproducing kernel Hilbert space where $g$ is a kernel $k\in\mathcal{H}$ and the supremum in (30) is thus attained. Such MMD-based optimizes only over the generator’s parameters but does not produce expressive generators, mainly because of the restriction imposed by the kernel structure. The second, instead, requires to learn both the generator and the discriminator, the latter in order to reach the supremum in (32). However, enforcing the Lipschitz constraint is not trivial and the alternation between generator and discriminator training suffers from the usual instability and slow convergence problems [5]. Even if the copula-based approach does not claim optimality, it possesses two desirable properties: compared to kernel-based methods, it uses a more powerful and appropriate discriminator, the copula density itself. Moreover, the fact that $c_{U}$ is obtained from a prior analysis renders the generator learning process uncoupled from the discriminator’s one.